If TC = 0.5q3 − 3q2 + 25q + 20 derive functions for: (a) MC, (b) AC, (c) the slope of AC.


1
Expert's answer
2021-08-02T10:50:49-0400
"MC=1,5q^3-6q+25"

"AC=0.5q^2-3q+25+\frac{20}{q}"

"\frac {\delta AC}{\delta q}=q-3-\frac{20}{q^2}"